OpenStudy (anonymous):

Find the value of sqrt(8/9)

OpenStudy (anonymous):

Anyone?

OpenStudy (ipwnbunnies):

Calculator?

OpenStudy (ipwnbunnies):

We can simplify it if that's what you're supposed to do....\[\frac{ \sqrt{8} }{ \sqrt{9} }\]

OpenStudy (ipwnbunnies):

= (2*sqrt(2))/3